# Robot Data Curation with Mutual Information Estimators

### v2

@@ Abstract
The performance of imitation learning policies often hinges on the datasets with which they are trained. Consequently, investment in data collection for robotics has grown across both industrial and academic labs. However, despite the marked increase in the quantity of demonstrations collected, little work has sought to assess the quality of said data despite mounting evidence of its importance in other areas such as vision and language. In this work, we take a critical step towards addressing the data quality in robotics. Given a dataset of demonstrations, we aim to estimate the relative quality of individual demonstrations in terms of both state diversity and action predictability. To do so, we estimate the average contribution of a trajectory towards the mutual information between states and actions in the entire dataset, which precisely captures both the entropy of the state distribution and the state-conditioned entropy of actions. Though commonly used mutual information estimators require vast amounts of data often beyond the scale available in robotics, we introduce a novel technique based on $k$-nearest neighbor estimates of mutual information on top of simple VAE embeddings of states and actions. Empirically, we demonstrate that our approach is able to partition demonstration datasets by quality according to human expert scores across a diverse set of benchmarks spanning simulation and real world environments. Moreover, training policies based on data filtered by our method leads to a 5-10% improvement in RoboMimic and better performance on real ALOHA and Franka setups. Work done as a Student Researcher at Google DeepMind Equal advising. Videos and code at https://jhejna.github.io/demonstration-info.
@@ I Introduction
Even though data quality is critical to the performance of imitation learning algorithms, little work has sought to measure it [6] and unfortunately, techniques used for data curation in vision and language do not transfer well to robotics. For example, n-gram classifiers have been extremely effective for web text [2] but are unable to handle high-dimensional continuous states and actions. Pre-trained models have been used to curate vision datasets [64], but are incapable of reasoning about actions. In contrast, we believe metrics for imitation learning should be able to measure both the diversity of states, which has been correlated with generalization performance [26, 43], and the relative quality of actions, which affects how well a policy is able to fit the expert distribution [6]. In this work we explore how both of these desiderata can be captured by the mutual information between states and actions.
Mutual information, or the bits of information learned about one random variable by observing another, precisely measures the difference between the marginal entropy of one variable and the conditional entropy of another. In the context of states and actions, this means that high mutual information encourages a large diversity of states (state entropy) but a predictable action distribution (low entropy of actions given a particular state). We thus propose using mutual information as a desirable metric for measuring data quality in imitation learning.
@@
When applying DemInf to a number of different robot platforms and environments, we find that it is able to consistently partition high- and low-quality data as scored by expert human annotators, outperforming both contemporary baselines and alternative mutual information estimators like InfoNCE [56]. Furthermore, using DemInf to subsample demonstration data, we are able to attain higher performing imitation learning policies across the RoboMimic [50] benchmark and real ALOHA policies on RoboCrowd Mirchandani et al. [52].
@@ III-A Imitation Learning
We consider the general problem of shaping the empirical expert distribution $\hat{\rho}_{\pi_{E}}(s)$ tabula rasa at the demonstration level. Specifically, our goal is to determine a score function $S(\tau)$ in a purely offline fashion that is able to predict the quality of demonstrations, where quality is determined by the performance of a policy trained with behavior cloning on the score-filtered demonstration dataset
@@ IV Mutual Information as a Quality Metric
$I(S;A)=H(S)-H(A\mid S)$ (4)
where $S$ and $A$ represent random variables for the state and action. First, we will discuss why minimizing the conditional action entropy allows for more accurate policies. Second, we discuss why maximizing state entropy can specifically improve performance in the sequential decision making setting.
@@
IV-A Minimizing Action Entropy
@@
IV-B Maximizing State Entropy
In addition to minimizing conditional action entropy, mutual information encourages high entropy in the state marginal distribution $H(S)$ (the first term of Eq. 4). Assuming we are unable to perfectly match the expert, during evaluation the learned policy $\pi_{\theta}$ is likely to make an imperfect prediction at a given state $s_{t}$ leading to a new state $s_{t+1}$. If this state is not close to any other state in the training dataset, the policy is unlikely to recover. However, if the dataset $\mathcal{D}_{N}$ contains higher coverage over the state space, we might expect to have a state similar to $s_{t+1}$ in the dataset.
Though mutual information is perhaps a natural metric for data curation, it can be practically difficult to estimate [19]. In the next section, we detail our approach to obtaining accurate estimates.
@@ V Method
In this section we propose the Demonstration Information Estimation (DemInf) method for computationally estimating mutual information. Though mutual information is usually considered at the distribution or dataset level, we are interested in scoring individual demonstrations for data curation. Thus, we measure the contribution of individual episodes to the overall mutual information of the dataset. Fortunately, this can easily be done as the majority of of empirical mutual information estimators can be decomposed into an average of sample-wise estimators.
@@ VI-A2 Baselines
Compatibility. Following Gandhi et al. [25], we use a measure of demonstration “compatibility” to score data. Namely, a demonstration is compatible with respect to a policy if it has either high “novelty” as measured by the prediction variance of an ensemble, or low novelty and high likelihood as measured by the average loss. In some sense, this could be akin to mutual information if novelty captures $H(S)$ and likelihood captures $H(A|S)$. Though this method was originally designed to be used in the online setting, we adopt it to the offline setting by training a policy on all data, then estimating the “compatibility” for each demonstration with respect to the overall policy.
@@ VI-A3 Architectures
For all state-based experiments we use MLPs with two hidden layers of size 512. For image-based experiments we use ResNet-18 Encoders [29] with spatial softmax activations following Mandlekar et al. [50], which are concatenated with state information as input to a MLP with two hidden layers of size 1024. When training VAEs from images we use matching ResNet-18 Decoder networks for each view. For each dataset we use the same architecture for all methods, where the latent $z$ dimension is set to be consistent across both DemInf and baselines. For all experiments we use the Adam optimizer with learning rate 0.0001 and a batch size of 256. State-based models are trained for 50,000 steps and image based models are trained for 100,000 steps using VMs provided by a Google TPU Research Cloud Grant. We run three seeds for all methods. More details and hyper-parameters can be found in Section -E.
@@ VI-C Mutual Information Estimators
Figure 8: Performance of ResNet18 + MLP BC Policies trained on filtered subsets of RoboMimic from Images. Evaluations are averaged over 200 trials for each of 3 seeds (600 total) after 100K training steps. Each dataset begins with 300 demonstrations.
@@ VI-D Does demonstration curation affect policy performance?
While comparing to ground truth labels allows us to assess the quality of different approaches to filtering, we ultimately care about the performance of downstream BC policies. In Fig. 8 we train BC policies on RoboMimic Can MH and Square MH from images when filtering different numbers of episodes from the dataset $\mathcal{D}_{N}$ according to the best baselines for the tasks. Overall, we find that at all data scales DemInf performs better than baselines, which exhibit far less consistent performance trends overall. For example, uncertainty often shows little improvement until the majority of the dataset is filtered. Crucially, we see that filtering data with DemInf performs better than training on all of the data by over 10% in Can and is the only method to improve upon training on all of the data in Square.
Figure 9: Performance of ACT trained on filtered versus non-filtered RoboCrowd (HiChew and TootsieRoll) and Diffusion Policy trained on filtered versus non-filtered data for the Franka tasks (DishRack and PenInCup). Evaluations for RoboCrowd are shown using the same scoring procedure from Mirchandani et al. [52] over 10 trials, and evaluations for Franka are completed over 15 trials.
This trend continues in real world evals for both RoboCrowd and Franka where we compare training policies on all of the demonstrations (no filter), training policies on a random 50% subset of the demonstrations, and filtering 50% of the demonstrations with DemInf in Fig. 9. We train ACT [76] for RoboCrowd and Diffusion Policy [14] for Franka. In RoboCrowd, after scoring trials according to the methodology in Mirchandani et al. [52] (1 for grasping any number of candies, 2 for returning it to the user, and 3 for returning only one as in the demonstrations), we find that the DemInf policy not only more commonly successfully completes the task, but also exhibits better motion when compared to training on all of the data. When considering the same number of demonstrations randomly selected from the dataset, we find that the gap in score is larger, indicating that better performance can be attained by collecting only good data. On the Franka tasks DemInf always outperforms the random subset, but training on all data does slightly better than filtering with DemInf on “PenInCup”. This is likely because “PenInCup”, being the harder task, requires more data for representation learning even if the quality is poor. However, when dataset size is kept at parity (50% random subset) we see a huge difference in performance, indicating that in all tasks data quality matters.
@@ VII-A Limitations
Action Entropy. DemInf considers the mutual information between states and actions to curate demonstrations. However, this ignores the fact that in the sequential setting, action entropy directly determines state entropy through time, e.g. a more random policy will visit more states. In this sense, DemInf may be over-indexing on $H(A|S)$. A better approach might be to consider measuring $I(S_{1};A_{1},...A_{T})$, or the mutual information between initial states and action sequences. However, doing so reduces the amount of data available for estimation by a factor of $T$, which is typically 100-1000, making this approach more challenging in practice.
State Entropy. Because of the $H(S)$ component, DemInf could prefer outlier trajectories that visit a new part of the state space, so long as there are no nearby trajectories with conflicting actions. For this reason as well, the aforementioned approach of measuring $I(S_{1};A_{1},...A_{T})$ might be preferred in a world of infinite data.
Pauses. Because DemInf considers the average estimated $\hat{I}$ across a trajectory, it is susceptible to preferring data that is predictable, but might not make progress towards completing the task. For example, if a robot pauses for an extended amount of time, the action distribution is very predictable. However, this behavior is not desired in practice. To mitigate this effect, we recommend ensuring that all data completes the task and pauses are filtered.

### v3

@@ Abstract
The performance of imitation learning policies often hinges on the datasets with which they are trained. Consequently, investment in data collection for robotics has grown across both industrial and academic labs. However, despite the marked increase in the quantity of demonstrations collected, little work has sought to assess the quality of said data despite mounting evidence of its importance in other areas such as vision and language. In this work, we take a critical step towards addressing the data quality in robotics. Given a dataset of demonstrations, we aim to estimate the relative quality of individual demonstrations in terms of both action diversity and predictability. To do so, we estimate the average contribution of a trajectory towards the mutual information between states and actions in the entire dataset, which captures both the entropy of the marginal action distribution and the state-conditioned action entropy. Though commonly used mutual information estimators require vast amounts of data often beyond the scale available in robotics, we introduce a novel technique based on $k$-nearest neighbor estimates of mutual information on top of simple VAE embeddings of states and actions. Empirically, we demonstrate that our approach is able to partition demonstration datasets by quality according to human expert scores across a diverse set of benchmarks spanning simulation and real world environments. Moreover, training policies based on data filtered by our method leads to a 5-10% improvement in RoboMimic and better performance on real ALOHA and Franka setups. Work done as a Student Researcher at Google DeepMind Equal advising. Videos and code at https://jhejna.github.io/demonstration-info.
@@ I Introduction
Even though data quality is critical to the performance of imitation learning algorithms, little work has sought to measure it [6] and unfortunately, techniques used for data curation in vision and language do not transfer well to robotics. For example, n-gram classifiers have been extremely effective for web text [2] but are unable to handle high-dimensional continuous states and actions. Pre-trained models have been used to curate vision datasets [64], but are incapable of reasoning about actions. In contrast, we believe metrics for imitation learning should be able to measure the relative predictability of the state-action distribution directly, which affects how well a policy is able to fit the expert [6]. In this work we explore how this desiderata can be captured by the mutual information between states and actions.
Mutual information, or the bits of information learned about one random variable by observing another, precisely measures the difference between the marginal entropy of one variable and the conditional entropy of another. In the context of states and actions, this means that high mutual information encourages a large diversity of actions (action entropy) but a predictable action distribution (low entropy of actions given a particular state). We thus propose using mutual information as a desirable metric for measuring data quality in imitation learning.
@@
When applying DemInf to a number of different robot platforms and environments, we find that it is able to consistently partition high- and low-quality data as scored by expert human annotators, outperforming both contemporary baselines and alternative mutual information estimators like InfoNCE [56]. Furthermore, using DemInf to subsample demonstration data, we are able to attain higher performing imitation learning policies across the RoboMimic [50] benchmark and real ALOHA policies on RoboCrowd [52].
@@ III-B Demonstration Curation
We consider the general problem of shaping the empirical expert distribution $\hat{\rho}_{\pi_{E}}(s)$ tabula rasa at the demonstration level, assuming all demonstrations are successful at completing the desired task. Specifically, our goal is to determine a score function $S(\tau)$ in a purely offline fashion that is able to predict the quality of demonstrations, where quality is determined by the performance of a policy trained with behavior cloning on the score-filtered demonstration dataset
@@ IV Mutual Information as a Quality Metric
$I(S;A)=H(A)-H(A\mid S)$ (4)
where $S$ and $A$ represent random variables for the state and action. First, we will discuss why minimizing the conditional action entropy allows for more accurate policies. Second, we discuss why maximizing the marginal action entropy is required to learn a good policy.
@@
IV-A Minimizing Conditional Action Entropy
@@
IV-B Maximizing Marginal Action Entropy
In addition to minimizing conditional action entropy, mutual information encourages high entropy in the marginal action distribution $H(A)$ (the first term of Eq. 4). While this might be puzzling at first, it has an important regularizing affect. Without considering the marginal entropy of actions $H(A)$, the conditional entropy $H(A|S)$ could be trivially minimized by distributions that have constant actions, e.g. taking the same action regardless of the state. Having a high marginal action entropy avoids this pitfall, forcing the learned policy to pay attention to the state when making predictions, which is desirable for closed-loop control.
We find mutual information to be a useful metric for data quality. However, practitioners in robotics have often considered simpler metrics as heuristics for data quality. For example, is often chosen as a convenient proxy for enabling generalization. Mutual information objective $I(S;A)$ can equivalently be decomposed as $H(S)-H(S|A)$ to uncover state entropy maximization. However, the latter term is less interpretable in this objective. As we will later average mutual information estimates across entire trajectories, DemInf will not explicitly optimize for initial state diversity, but instead the underlying predictability of the data.
@@ V Method
Though mutual information is perhaps a natural metric for data curation, it can be practically difficult to estimate [19]. In this section we propose the Demonstration Information Estimation (DemInf) method for computationally estimating mutual information for demonstration data. Though mutual information is usually considered at the distribution or dataset level, we are interested in scoring individual demonstrations for data curation. Thus, we measure the contribution of individual episodes to the overall mutual information of the dataset. Fortunately, this can easily be done as the majority of of empirical mutual information estimators can be decomposed into an average of sample-wise estimators.
@@ VI-A2 Baselines
Compatibility. Following Gandhi et al. [25], we use a measure of demonstration “compatibility” to score data. Namely, a demonstration is compatible with respect to a policy if it has either high “novelty” as measured by the prediction variance of an ensemble, or low novelty and high likelihood as measured by the average loss. Though this method was originally designed to be used in the online setting, we adopt it to the offline setting by training a policy on all data, then estimating the “compatibility” for each demonstration with respect to the overall policy.
@@ VI-A3 Architectures
For all state-based experiments we use MLPs with two hidden layers of size 512. For image-based experiments we use ResNet-18 Encoders [29] with spatial softmax activations following Mandlekar et al. [50], which are concatenated with state information as input to a MLP with two hidden layers of size 1024. When training VAEs from images we use matching ResNet-18 Decoder networks for each view. For each dataset we use the same architecture for all methods, where the latent $z$ dimension is set to be consistent across both DemInf and baselines. Images resized to $84\times 84$ for RoboMimic and $128\times 128$ otherwise. For all experiments we use the Adam optimizer with learning rate 0.0001 and a batch size of 256. State-based models are trained for 50,000 steps and image based models are trained for 100,000 steps using VMs provided by a Google TPU Research Cloud Grant. We run three seeds for all methods. More details and hyper-parameters can be found in Section -E.
@@ VI-C Mutual Information Estimators
Figure 8: Performance of ResNet18 + MLP BC Policies trained on filtered subsets of RoboMimic from Images. Evaluations are averaged over 200 trials for each of 3 seeds (600 total) after 100K training steps. Each dataset begins with 300 demonstrations, except for “Square Mh 1000” which has 1000.
Figure 9: Performance of ACT trained on filtered versus non-filtered RoboCrowd (HiChew and TootsieRoll) and Diffusion Policy trained on filtered versus non-filtered data for the Franka tasks (DishRack and PenInCup). Evaluations for RoboCrowd are shown using the same scoring procedure from Mirchandani et al. [52] over 10 trials, and evaluations for Franka are computed using binary success over 15 trials.
@@ VI-D Does demonstration curation affect policy performance?
While comparing to ground truth labels allows us to assess the quality of different approaches to filtering, we ultimately care about the performance of downstream BC policies. In Fig. 8 we train BC policies on RoboMimic Can MH and Square MH from images when filtering different numbers of episodes from the dataset $\mathcal{D}_{N}$ according to the best baselines for the tasks. Overall, we find that at all data scales DemInf performs better than baselines, which exhibit far less consistent performance trends overall. For example, uncertainty often shows little improvement until the majority of the dataset is filtered. Crucially, we see that filtering data with DemInf performs better than training on all of the data by over 10% in Can and is the only method to improve upon training on all of the data in Square. To further assess performance on larger datasets, we add an additional 700 successful rollouts from Diffusion Policy [14] to Square MH to bring the total to 1000 demos for “Square MH 1000”. Even with larger datasets, DemInf maintains performance. We include results for robomimic state in Section -C.
This trend continues in real world evals for both RoboCrowd and Franka where we compare training policies on all of the demonstrations (no filter), training policies on a random 50% subset of the demonstrations, and filtering 50% of the demonstrations with DemInf in Fig. 9. We train ACT [76] for RoboCrowd and Diffusion Policy [14] for Franka. In RoboCrowd, after scoring trials according to the methodology in Mirchandani et al. [52] (1 for grasping any number of candies, 2 for returning it to the user, and 3 for returning only one as in the demonstrations), we find that the DemInf policy not only more commonly successfully completes the task, but also exhibits better motion when compared to training on all of the data. When considering the same number of demonstrations randomly selected from the dataset, we find that the gap in score is larger, indicating that better performance can be attained by collecting only good data. On the Franka tasks DemInf always outperforms the random subset, but training on all data does slightly better than filtering with DemInf on “PenInCup”. This is likely because “PenInCup”, being the harder task, requires more data for representation learning even if the quality is poor. However, when dataset size is kept at parity (50% random subset) we see a huge difference in performance, indicating that in all tasks data quality matters. VIP does well on the Franka tasks as well, which can be attributed to the fact that the purposefully suboptimal demonstrations we collected were significantly longer than the expert demonstrations, making filtering via a value function particularly effect.
@@ VII-A Limitations
Time Evolution. DemInf considers the aggregate mutual information between states and actions within the dataset. Our analyses, however, largely ignores the fact that these quantities are linked in the sequential setting via the environment dynamics. Concretely, we can increase the marginal action entropy with a more random policy, which will likely increase the state entropy as well through the dynamics. However, this comes at the cost of increasing $H(A|S)$, and it is unclear how DemInf balances these factors. An approach to disentangle the effects of dynamics might consider measuring $I(S_{1};A_{1},\dots,A_{T})$, or the mutual information between initial states and action sequences. However, doing so reduces the amount of data available for estimation by a factor of $T$, which is typically 100-1000, making this approach more challenging in practice.
Pauses. Because DemInf considers the average estimated $\hat{I}$ across a trajectory, it is susceptible to preferring data that is predictable, but might not make progress towards completing the task. For example, if a robot pauses for an extended amount of time at a particular state, the action distribution is very predictable. However, this behavior is not desired in practice. To mitigate this effect, we recommend ensuring that all data completes the task and pauses are filtered.
@@
Success. DemInf assumes that all provided demonstrations are successful at completing the desired task.